Two charges 2C and -2C are at 0.01m. Calculate the electric field intensity 0.3m in equatorial plane.


1
Expert's answer
2021-06-07T08:30:35-0400


Let us consider two equal and opposite charges "q" and "-q" at points A and B separated by a distance "2d" as shown in the figure. We have to find the electric field at an equatorial point "P" distant "x" from the center "O" .

From "\Delta AOP" and "\Delta BOP" ,

"AP=BP=\sqrt{x^2+d^2}"

The electric field at P due to "q" is

"\vec{E}_a=\frac{1}{4\pi \epsilon_0}\frac{q}{x^2+d^2}" directed along AP.

The electric field at P due to "-q" is

"\vec{E}_b=\frac{1}{4\pi \epsilon_0}\frac{q}{x^2+d^2}" directed along PB.

Note that the electric fields "\vec{E}_a" and "\vec{E}_b" have equal magnitude i.e,

"|\vec{E}_a|=|\vec{E}_b|=E(say)"

To find the resultant electric field at "P" we will resolve "\vec{E}_a" and "\vec{E}_b" in horizontal and vertical components as shown below.



The vertical components "E_a\sin\theta" and "E_b\sin\theta" will cancel each other (since "|\vec{E}_a|=|\vec{E}_b|" ).

The horizontal components "E_a\cos\theta" and "E_b\cos\theta" will add up to give the resultant electric field.

"\therefore" Resultant electric field at "P" is

"E_r=E_a\cos\theta+E_b\cos\theta\\
\Rightarrow E_r=2E\cos\theta\\
\Rightarrow E_r=2\frac{1}{4\pi\epsilon_0}\frac{q}{x^2+d^2}.\frac{d}{\sqrt{x^2+d^2}}\\
\Rightarrow \boxed{ E_r=\frac{1}{4\pi\epsilon_0}\frac{q.2d}{(x^2+d^2)^{3/2}}}" directed along AB.

Given, "q=2C, 2d=0.01\ m, x=0.3\ m"

We know "\frac{1}{4\pi\epsilon_0}=9\times 10^9 N.m^2.C^{-2}"

"\therefore E_r=9\times 10^9\times \frac{2\times 0.01}{(0.3^2+0.005^2)^{3/2}}\ N/C\\
 E_r=6.7\times 10^9\ N/C"


Answer: The electric filed intensity at 0.3m in equatorial plane is "6.7\times 10^9 \ N/C"